• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 05:40:58 on June22,2004

    Entry number 126621

    This entry is a followup to: 126582

    Followups:

    keyword=Current ramp study details (late post)

    h2>Target "boiling" septum heating study - June 21 2004
       D. Armstrong, K. Kumar, H. Benaoum, K. Aniol, et al.
    
     CHL flow = 22.2 g/s, JT valver = 63%, 
     target input temps: 7.4 K, output temp 8.2 K
    
    - raster = 3.5 x 3.5 (unless otherwise noted)
    - at each setting, let septa temps stabilize so that stable to within +-0.05 K for 5 minutes
                                                   Septa
    Run   Event Beam        target             Right     Left
            #            fan  HPHtr Press  upper lower upper lower 
           (K)  (muA)   (Hz)  (W)  (psi)    (K)   (K)  (K)    (K)
    --------------------------------------------------------------------------
    2335         4.5     48     -    221    6.70  6.40  6.15  6.30
    2336         9.4     48    443   219    6.85  6.55  6.30  6.45
    2337        18.9     48    403   220    7.10  6.80  6.50  6.65
    2344-2346   28.8     48    385   220    7.55  7.30  6.80  6.95
    2347         "       48     "     "      "      "    "     "    (1x1 raster)
    2348        33.5     48    353   220    7.80  7.60  6.90  7.10
    2349        37.7     48    325   221    7.80  7.60  7.00  7.20
         Right Septum trips at these temperatures.
    
    2350        44.3     48    305   221     -     -    7.20  7.40
    2351         "       "      "     "      -     -      "    "
    2352        48.3     48    274   221     -     -      "    " 
     "   12.5    "       60    225   219     -     -    7.25  7.45                      
     "   18.2    "       70    192   219     -     -    7.30  7.50
     "   24.0    "       24    364   231     -     -    7.30  7.50
    2353         "       48     "     "      -     -      "    "   (oversample =13)
    2354        52.1     48    260   221     -     -    7.45  7.65
          9      "       60    198   219     -     -      "    "
         14.5    "       70    168   219     -     -      "    "
    2355        57.3     70    139   218     -     -    7.50  7.70
          6.3    "       48    226   217     -     -      "    "
         19.5    "       24    337   232     -     -    Quenches
            Left septum quenches at these temperatures
    
      try for higher currents, but Hall A laser attenuator maxed out. 
         OUCH!!
    
    2356        52.9     48    240   220     -    -      -     -
    
      try to increase raster size. MCC nis at limit, which somehow did not
    get increase, despite Bob's efforts. End the study, back to production.
    
     Conclusions:
    
    1) Right septum quenches (again, very reproducible) at 7.80 K on upper coil,
      which is at about 34 µA on Helium cell
    
    2) Left septum quenches at 7.5 K (upper coil) 7.7 K (lower coil), which
       is at about 57 µA on Helium cell
    
    3) Target had lots of cooling headroom above 57 µA; depending on fan speed, at this operating temperature, we could get to 95 µA (48 Hz fan speed) using 22.5 g/s of 4.5 K coolant from CHL; with a better fan which dumps less heat we could get higher.
    
    4) Raster size (again) doies not seem to affect septum heating, at least at 29 µA. 
    
    5) Offline analysis need to look at "boiling" at the higehr currents, but we got various fan speeds at 48, 52 and 57 µA, so should be enough data to see how things are at these currents. 
    
    6) Too bad we could not get to higher currents because of reduced QE on the photocathode.